# Gender differences in collaboration and career progression in physics

### v1

@@ Abstract
We examine gender differences in collaboration networks and academic career progression in physics. We use the likelihood and time to become a principal investigator (PI) and the length of an author’s career to measure career progression. Utilising logistic regression and accelerated failure time models, we examine whether the effect of collaboration behaviour varies by gender. We find that, controlling for the number of publications, the relationship between collaborative behaviour and career progression is almost the same for men and women. Specifically, we find that those who eventually reach principal investigator (PI) status, tend to have published with more unique collaborators. In contrast, publishing repeatedly with the same highly interconnected collaborators and/or larger number of coauthors per publication is characteristic of shorter career lengths and those that do not attain PI status. We observe that the tie strength is stronger for women than men, and women tend to collaborate in more tightly connected and larger groups than men. Finally, we observe that women are less likely to attain the status of PI throughout their careers and have a lower survival probability compared to men, which calls for policies to close this crucial gap.
@@ 1 Introduction
The field of physics has a long history of gender disparities, in which the majority of researchers have been men. Research has shown that women are underrepresented at all levels of physics education and in the physics workforce [1, 2]. Furthermore, female researchers in physics have been found to encounter greater barriers to professional advancement and tend to leave academia earlier than their male counterparts [3, 4, 5]. A 2014 report indicates that women represent 23% of assistant professors, 18% of associate professors, and only 10% of full professors in physics departments in the United States [6]. Similar findings apply to Europe and other areas of STEM fields [7]. But what are the driving factors that help or hinder women from the career progression in physics? Here, we examine that question through the association between gender, collaboration behaviour and career progression.
@@
In recent years, the impact of collaboration on career progression has garnered increased attention as a potential contributing factor to these disparities [8]. Collaboration is a crucial aspect of scientific research and has been shown to significantly impact professional development and advancement [9, 10]. A number of studies have shown that women have smaller collaboration networks than men [11, 12, 13, 14]. This difference in collaboration behaviour may negatively impact the career advancement of women, as collaborations are often required for scientific success and can lead to greater visibility and credibility in the field [15].
@@
Furthermore, women are underrepresented in leadership positions in academia and industry. For example, Ginther and Kahn  [16] found that women in scientific, engineering, and medical fields are less likely to be promoted to full professor than men, and similarly, Moss-Racusin et al. [17] found that women scientists are less likely to be hired and promoted. These findings indicate a persistent disparity in the progression of women towards leadership roles in academia and industry. Career progression is not only important for individuals themselves, but can also affect the structure of the professional hierarchy for future generations. For example, more diversity and inclusion on senior managements levels has been seen to affect policies and research priorities [18]. Besides, peer mentoring with individuals of the same gender during crucial developmental stages enhances the success and retention of women, producing long-term benefits, especially in more male-dominated fields such as science and engineering [19, 20].
@@
In this paper, we present a comprehensive analysis of publication data from the American Physical Society (APS) that reveals several key findings (Figure 1). We find differences in the way men and women collaborate. Women have stronger connections with each coauthor, collaborators who are more closely connected to each other and a greater mean number coauthors per publication. We identify differences in the career progression of men and women, finding that women exhibit lower longevity than their male counterparts regardless of whether they attain a PI status. This indicates that the career progression of females in academia may not fully capture the value of these dense and collaborative networks.
@@
Figure 1: Illustration of career progression and collaboration. (A) Author $a$ collaborates with other scientists ($b$, $c$, $d$ and $e$) to publish seven articles over their career (solid circles). Upon their third last-author publication (green border), $a$ achieves Principal Investigator (PI) status (green box). For all scientists who eventually become PI, we calculate the time to become PI as the number of years between their first publication and the publication through which they attained PI status. The number of years after becoming PI until the last publication define the time after PI. Using the careers of all focal authors, we further compute career lengths as the time passed between their first and last published articles and the likelihood of becoming a PI. (B) To measure collaboration behaviour, we aggregate authors’ ego networks over their careers. In this example, the focal author $a$ has a network size of four unique coauthors from seven publications. She has a mean tie strength of $(11/4=2.75)$, local clustering coefficient of $(4/6=0.67)$, and $(11/7=1.57)$ coauthors on average per publication.
@@ 2 Selecting focal authors
We analyse a dataset from the American Physical Society (APS) between 1893 and 2020, comprising $678,916$ publications in 19 different Physical Review journals (further details given in table S1). Within this dataset, we use an algorithm for name disambiguation based on the method used by Sinatra et al. [21] to identify unique authors and employ a comprehensive methodology [22] to infer the gender of each author (see supplementary  S1 information for full details). Using this process we identify the gender of $230,086$ unique authors.
@@
    $157,460$ authors are excluded due to having published fewer than three papers or more than $400$ papers in total.
@@
    $11,051$ authors are excluded due to an interval year between first publication and last publication that is smaller than $3$ years or larger than $50$ years.
@@
    $541$ authors are excluded due to having less than two unique coauthors.
@@
    $37,707$ active authors are excluded because they are still publishing after 2015.
After applying these filtering criteria, $23,327$ focal authors remain in the dataset, including $2,331$ women and $20,996$ men.
@@ 3 Collaboration behaviour
To capture collaboration behaviour, we construct a weighted egocentric network for each author, derived from joint publications with other authors (Figure 1). In each egocentric network, the ego $E$, represents the focal author and alters are coauthors who have published papers with the focal author. The edge weight, or tie strength, indicates the number of coauthored publications. An edge, or connection, exists between a pair of alters if they have coauthored at least one paper together with the ego. This choice of representation concentrates on the focal author’s direct collaborations at the cost of omitting some connections between alters that do not relate to the focal author. We calculated four statistics for each egocentric network to measure the collaboration behaviour of the focal authors (See Figure 1 for details):
@@
  • Local clustering coefficient is the proportion of pairs of alters that are connected to each other in the network. The clustering coefficient indicates the degree to which the coauthors are interconnected among themselves.
  • Coauthors per publication is the sum of the edge weights in the network divided by the total number of publications authored by the ego. The mean number of coauthors per publication indicates the size of the collaborations that the ego tends to engage in.
@@ 3.1 Adjusting for differences in number of publications
The number of papers a focal author publishes can have an influence on the pattern of collaborations that the focal author engages in. For instance, the more a focal author publishes, the more opportunity the focal author has to increase their network size. We examine the effect of this potential confounder by setting up a linear regression to estimate an author’s number of publications according to their collaboration statistics. Table 1 displays the regression coefficients that suggest a significant association between publication count and collaboration statistics. To address this confounding factor and allow us to more closely examine the relationship between gender and collaboration behaviour, we make an adjustment to the collaboration statistics. Specifically, we use model residuals to capture the collaboration behaviour with the effects of the publication number removed (further details given in supplementary material  S2)  [23]. We denote these statistics that account for the potential influence of publication count as “adjusted”.
@@
network size tie strength clustering coefficient coauthors per publication
@@ 3.2 Correlation between collaboration statistics
All pairs of collaboration statistics are found to be statistically significant, however, for most pairs, the effect size is relatively small as the estimated coefficients are close to zero. Exceptions are network size with coauthors per publication and tie strength with clustering coefficient. Since these statistics are adjusted for the number of publications, which remove the effect of publication count, it is expected that the network size would correlate with the mean number of coauthors per publication. When an author has a larger network size, they tend to have more coauthors per paper. The moderate correlation between tie strength and clustering coefficient seems to indicate that repeated co-authorship is associated with triadic closure, i.e., coauthors become more interconnected. We also see that tie strength and clustering coefficient have a weak negative correlation with network size. This relationship is likely due to the fact that as the network size increases, the collaboration network can become less dense and repeated collaborations occur less frequently, after adjusting for the number of publications.
@@ 3.3 Gender differences in collaboration behaviour
In order to investigate the difference in collaborative behaviour between male and female authors, we examine both the distributions of collaborative behaviour and the mean collaboration statistics derived from Generalised Linear Model (GLM) regression models. This dual approach helps us to discern potential disparities in collaboration patterns between male and female authors.
Figure 3 presents the density distributions of adjusted collaborative behaviour statistics for both genders, normalised based on the number of publications. These visualisations reveal that the distributions of network size are indistinguishable across genders when accounting for the number of published papers. On the contrary, female authors seem to have greater mean tie strength, clustering coefficient and average number of coauthors per publication compared to male authors. GLM coefficients (S2) confirm these differences in means to be statistically significant.
Figure 3: Gender differences in collaboration statistics. We show the distribution of collaboration statistics. Blue curves represent male scientists, and orange curves represent female scientists. Vertical lines indicate the mean value.
@@ 3.4 Assortative mixing patterns in collaborations
Here we investigate to what extent collaboration behaviour is similar between collaborators. For instance, focal authors who frequently collaborate together might also start to collaborate in similar ways. We use network assortativity [24] to assess the extent to which focal authors tend to exhibit similar collaboration characteristics as their coauthors. We stratify the author pairs into distinct collaboration tiers based on the number of papers they coauthored. At each tier, we calculate the assortativity of the collaboration statistics between coauthors to determine whether or not the frequency of collaboration effects the level of assortativity. Note that assortativity values are based on the residual collaboration statistics once we remove the author pair under consideration. For example, when calculating the assortativity of mean tie strength between a pair of authors $a$ and $b$, we calculate their mean tie strengths without including the tie between $a$ and $b$.
Figure 4 shows the generally positive assortativity for all four collaboration statistics, indicating that authors with similar collaboration styles tend to collaborate together. The tie-strength assortativity grows with the number of coauthored publications. Often repeated ties link pairs of scientists with similar average tie strengths in their local networks. In contrast, the assortativity of coauthors per publication decreases with repeated collaboration. Coauthors that publish only once or twice are most similar in terms of their average number of coauthors per publication.
Figure 4: Assortativity of collaboration statistics according to the number of co-authored publications. The heat map provides a visual representation of the assortativity of each of the collaboration statistics (y-axis) between co-author pairs. The bar chart along the top shows the logarithmic frequency of the co-authorship pairs.
@@ 4.1 Gender differences in career progression
To investigate gender differences in academic career progression, we examine three key aspects: the likelihood of becoming a PI, overall career length, and career length after becoming a PI. We use the cumulative distribution function to show the probability of becoming a PI for each gender by a given time. We visualise the overall career length and career length after PI using the complementary cumulative distribution function for male and female authors. We can assess whether there are profound disparities in career progression between male and female authors by comparing these curves.
@@
Figure 5: Gender differences in career progression. These graphs show cumulative probability of being a PI and complementary cumulative probability during career length and career length after PI. Solid coloured lines represent cumulative and complementary cumulative density functions and the shaded areas indicate the 95% confidence interval.
@@ 4.2 Assortativity in career progression
Figure 6 shows the assortativity of time spent at different career stages between coauthors. Overall we see slight positive assortativity in the career progression statistics. There does not appear to be any clear pattern in relation to the number of coauthored publications. These results suggest that frequent collaboration has little or no effect on the similarity of a co-author’s career progression.
Figure 6: Assortativity of career progression statistics according to the number of co-authored publications. The heat map provides a visual representation of the assortativity of each of the career progression statistics (y-axis) pairs of co-authors. The colour of each cell represents the assortativity according to the number of co-authored publications (y-axis). The bar chart along the top shows the logarithmic frequency of the co-authorship pairs.
@@ 5 Relationship between collaboration behaviour and career progression
Now we explore the relationship between collaboration behaviour and career progression, and whether or not this relationship varies by gender. We use logistic regression models to investigate the relationship between collaboration statistics and the likelihood of an individual becoming a Principal Investigator (PI) and Accelerated Failure Time Models (AFTs) to analyse the association between collaboration and the time spent at different career stages.
@@ 5.1 Likelihood of becoming a PI
We assess the relationship between each of the adjusted collaboration statistics (network size, tie strength, clustering coefficient, and coauthors per publication) and the likelihood of becoming a PI using logistic regression models. We also include gender and the interaction between gender and the collaboration statistic as predictors. We model the probability of an individual $i$ succeeding to become a PI ($s_{i}=1$) as follows:
@@
where $g_{i}$ is a binary variable that denotes the gender of the $i$-th individual, where $g_{i}=1$ indicates female and $g_{i}=0$ indicates male. The variable $c_{i}$ corresponds to the value of the specific collaboration statistic for the $i$-th individual. The model includes $\beta_{0}$ as the intercept term, $\beta_{1}$ for the collaboration static effect, $\beta_{2}$ for the gender effect, and $\beta_{3}$ for the interaction effect between gender and the collaboration statistic. We also construct two separate models, one for male and one for female focal authors, to analyse the relationship within the two gender groups by fixing $\beta_{2}=\beta_{3}=0$.
Figure 7 is a forest plot illustrating the association between the adjusted collaboration statistics and the likelihood of becoming a Principal Investigator (PI). The plot includes three types of model (interaction, female, and male) for each adjusted collaboration statistic (network size, tie strength, clustering coefficient, and coauthors per publication). Each row represents the value of the relevant model coefficient given in Eq. (1). For the interaction models we show the coefficient $\beta_{3}$ that captures the interaction effect between gender and the collaboration statistic. For the female models, the combined effect of the collaboration statistic is represented by the coefficient $\beta_{1}+\beta_{3}$. For the male models, the effect is represented by the coefficient $\beta_{1}$.
Figure 7 shows that the interaction terms between gender and the four collaboration statistics are not statistically significant in any of the models, suggesting no substantial moderating effect of gender on these relationships. We see that an increased network size is significantly associated with a higher likelihood of becoming a PI, while a lower clustering coefficient, number of coauthors per publication and mean tie strength are significantly associated with a greater likelihood of becoming a PI. However, the smaller number of female authors makes it harder to assess the statistical significance between mean tie strength and the likelihood of being a PI for women.
@@
Figure 7: The forest plot of adjusted collaboration statistics on likelihood of being a PI. We present three types of model for each type of collaboration statistic (network size, tie strength, clustering coefficient and coauthors per publication). The interaction model refers to the full model described in Eq. (1) and includes the interaction effect of gender with the collaboration statistic. The female and male models assess the effect of the respective collaboration statistic on the likelihood of being a principal investigator (PI) for the separate groups of female and male authors. The Number column shows the sample size of each model. The markers show the coefficient estimates, with error bars indicating the 95% confidence intervals (CIs) for each estimate.

### v2

@@ Abstract
We examine gender differences in collaboration networks and academic career progression in physics. We use the likelihood and time to become a principal investigator (PI) and the length of an author’s career to measure career progression. Utilising logistic regression and accelerated failure time models, we examine whether the effect of collaboration behaviour varies by gender. We find that, controlling for the number of publications, the relationship between collaborative behaviour and career progression is almost the same for men and women. Specifically, we find that those who eventually reach principal investigator (PI) status, tend to have published with more unique collaborators. In contrast, publishing repeatedly with the same highly interconnected collaborators and/or larger number of co-authors per publication is characteristic of shorter career lengths and not attaining PI status. We observe that women tend to collaborate in more tightly connected and larger groups than men. Finally, we observe that women are less likely to attain the status of PI throughout their careers and have a lower survival probability compared to men, which calls for policies to close this crucial gap.
@@ 1 Introduction
The field of physics has a long history of gender disparities, in which the majority of researchers have been men. Research has shown that women are underrepresented at all levels of physics education and in the physics workforce [1, 2]. Furthermore, female researchers in physics have been found to encounter greater barriers to professional advancement and tend to leave academia earlier than their male counterparts [3, 4, 5]. A 2014 report indicates that women represent 23% of assistant professors, 18% of associate professors and only 10% of full professors in physics departments in the United States [6]. Similar findings apply to Europe and other areas of STEM fields [7]. But what are the driving factors that help or hinder women from the career progression in physics? Here, we examine that question through the association between gender, collaboration behaviour and career progression.
@@
In recent years, the impact of collaboration on career progression has garnered increased attention as a potential contributing factor to these disparities [8]. Collaboration is a crucial aspect of scientific research and has been shown to significantly impact professional development and advancement [9, 10]. A number of studies have shown that women have smaller collaboration networks than men [11, 12, 13, 14]. This difference in collaboration behaviour may negatively impact the career advancement of women, as collaborations are often required for scientific success and can lead to greater visibility and credibility in the field [15]. Furthermore, Li et al. [16] found that differences in collaboration networks could largely explain gendered disparities in productivity and impact.
@@
Women are under-represented in leadership positions in academia and industry. Ginther and Kahn [17] found that women in scientific, engineering and medical fields are less likely to be promoted to full professor than men. Similarly, Moss-Racusin et al. [18] found that women scientists are less likely to be hired and promoted. These findings indicate a persistent disparity in the progression of women towards leadership roles in academia and industry. Career progression is not only important for individuals themselves, but can also affect the structure of the professional hierarchy for future generations. For example, more diversity and inclusion at senior management levels has been seen to affect policies and research priorities [19]. Besides, peer mentoring with individuals of the same gender during crucial developmental stages enhances the success and retention of women, producing long-term benefits, especially in more male-dominated fields such as science and engineering [20, 21].
@@
In this paper, we present a comprehensive analysis of publication data from the American Physical Society (APS) that reveals several key findings (Figure 1). We find differences in the way men and women collaborate. Women have stronger connections with each co-author, collaborators who are more closely connected to each other and a greater mean number co-authors per publication. We identify differences in the career progression of men and women, finding that women exhibit lower longevity than their male counterparts regardless of whether they attain a PI status. This indicates that the career progression of females in academia may not fully capture the value of these dense and collaborative networks.
@@
Figure 1: Illustration of career progression and collaboration. (A) The career trajectory of a focal author $a$, visualised as a timeline of publication events. Each solid black dot on the timeline represents a publication in time. Below each dot, the grey box illustrates the co-authorship network for a paper published at that time. Within each box, nodes represent co-authors; a green-outlined node indicates the last author on that publication. The solid green dot on the timeline marks the point at which $a$ becomes a Principal Investigator (PI), defined as having published their third last-author paper. The corresponding publication is highlighted with a green box. For all scientists who eventually become PI, we calculate the time to become PI as the number of years between their first publication and the publication through which they attained PI status. The number of years after becoming PI until the last publication define the time after PI. Using the careers of all focal authors, we further compute career lengths as the time passed between their first and last published papers and the likelihood of becoming a PI as the proportion of PIs among all authors. (B) The ego network of author $a$ aggregated over their entire career. Edges represent co-authorships with edge weights indicating number of shared papers. Here, author $a$ has collaborated with four unique co-authors across seven publications, yielding a network size of 4, mean tie strength of $11/4=2.75$, clustering coefficient of $4/6=0.67$ and an average of $11/7=1.57$ co-authors per paper.
@@ 2 Selecting focal authors
We analyse a dataset from the American Physical Society (APS) between 1893 and 2020, comprising $678,916$ publications in 19 different Physical Review journals (further details given in table S1). Within this dataset, we use an algorithm for name disambiguation based on the method used by Sinatra et al. [22] to identify unique authors and employ a comprehensive methodology [23] to infer the gender of each author (see supplementary S1 information for full details). Using this process we identify the gender of $248,788$ unique authors.
@@
    $191,300$ authors are excluded due to having published fewer than three papers or more than $400$ papers in total.
@@
    $8,708$ authors are excluded due to an interval year between first publication and last publication that is smaller than $3$ years or larger than $50$ years.
@@
    $464$ authors are excluded due to having less than two unique co-authors.
@@
    $29,217$ active authors are excluded because they are still publishing after 2015.
After applying these filtering criteria, $19,099$ focal authors remain in the dataset, who have collectively published $191,955$ papers, including $1,486$ women and $17,613$ men.
@@ 3 Collaboration behaviour
To capture collaboration behaviour, we construct a weighted egocentric network, for each author, derived from joint publications with other authors (Figure 1). In each egocentric network, the ego $E$ represents the focal author and alters are co-authors who have published papers with the focal author. The edge weight or tie strength, indicates the number of co-authored publications. An edge or connection, exists between a pair of alters if they have co-authored at least one paper together with the ego. This choice of representation concentrates on the focal author’s direct collaborations at the cost of omitting some connections between alters that do not relate to the focal author. We calculated four statistics for each egocentric network to measure the collaboration behaviour of the focal authors (see Figure 1 for details):
@@
  • Local clustering coefficient is the proportion of pairs of alters that are connected to each other in the network. The clustering coefficient indicates the degree to which the co-authors are interconnected among themselves. Throughout the paper, we describe ego networks with high clustering coefficients as “tight”, reflecting strong local interconnectedness among co-authors.
  • Co-authors per publication is the sum of the edge weights in the network divided by the total number of publications authored by the ego. The mean number of co-authors per publication indicates the size of the collaborations that the ego tends to engage in.
@@ 3.1 Adjusting for differences in number of publications
The number of papers a focal author publishes can have an influence on the pattern of collaborations that the focal author engages in. For instance, the more a focal author publishes, the more opportunity the focal author has to increase their network size. We examine the effect of this potential confounder by setting up a linear regression to estimate an author’s number of publications according to their collaboration statistics. Table 1 displays the regression coefficients that suggest a significant association between publication count and collaboration statistics. To address this confounding factor and allow us to more closely examine the relationship between gender and collaboration behaviour, we make an adjustment to the collaboration statistics. Specifically, we use model residuals to capture the collaboration behaviour with the effects of the publication number removed (further details given in supplementary material S2[24]. We denote these statistics that account for the potential influence of publication count as “adjusted”.
@@
network size tie strength clustering coefficient co-authors per publication
@@ 3.2 Correlation between collaboration statistics
All pairs of collaboration statistics are found to be statistically significant, however, for most pairs, the effect size is relatively small as the estimated coefficients are close to zero. Exceptions are network size with co-authors per publication and tie strength with clustering coefficient. Since these statistics are adjusted for the number of publications, which remove the effect of publication count, it is expected that the network size would correlate with the mean number of co-authors per publication. When an author has a larger network size, they tend to have more co-authors per paper. The moderate correlation between tie strength and clustering coefficient seems to indicate that repeated co-authorship is associated with triadic closure, i.e., co-authors become more interconnected. We also see that tie strength and clustering coefficient have a weak negative correlation with network size. This relationship is likely due to the fact that as the network size increases, the collaboration network can become less dense and repeated collaborations occur less frequently, after adjusting for the number of publications.
@@ 3.3 Gender differences in collaboration behaviour
In order to investigate the difference in collaborative behaviour between male and female authors, we examine both the distributions of collaborative behaviour and the mean collaboration statistics derived from Generalised Linear Models (GLMs). This dual approach helps us to discern potential disparities in collaboration patterns between male and female authors.
Figure 3 presents the density distributions of adjusted collaborative behaviour statistics for both genders, normalised based on the number of publications. These visualisations reveal that the distributions of network size are indistinguishable across genders when accounting for the number of published papers. On the contrary, female authors seem to have greater mean tie strength, clustering coefficient and average number of co-authors per publication compared to male authors. GLM coefficients (see supplementary S3 for details) confirm these differences in means to be statistically significant.
Figure 3: Gender differences in collaboration statistics. We show the distribution of collaboration statistics. Blue curves represent male scientists and orange curves represent female scientists. Vertical lines indicate the mean value.
@@ 3.4 Assortative mixing patterns in collaborations
Here we investigate to what extent collaboration behaviour is similar between collaborators. For instance, focal authors who frequently collaborate together might also start to collaborate in similar ways. We use network assortativity [25] to assess the extent to which focal authors tend to exhibit similar collaboration characteristics as their co-authors. We stratify the author pairs into distinct collaboration tiers based on the number of papers they co-authored. At each tier, we calculate the assortativity of the collaboration statistics between co-authors to determine whether or not the frequency of collaboration affects the level of assortativity. Note that assortativity values are based on the residual collaboration statistics once we remove the author pair under consideration. For example, when calculating the assortativity of mean tie strength between a pair of authors $a$ and $b$, we calculate their mean tie strengths without including the tie between $a$ and $b$.
Figure 4 shows the generally positive assortativity for all four collaboration statistics, indicating that authors with similar collaboration styles tend to collaborate together. The tie-strength assortativity grows with the number of co-authored publications. Often repeated ties link pairs of scientists with similar average tie strengths in their local networks. In contrast, the assortativity of co-authors per publication decreases with repeated collaboration. Co-authors that publish only once or twice are most similar in terms of their average number of co-authors per publication.
Figure 4: Assortativity of collaboration statistics according to the number of co-authored publications. The heat map provides a visual representation of the assortativity of each of the collaboration statistics (y-axis) between co-author pairs. The bar chart along the top shows the logarithmic frequency of the co-authorship pairs. A positive assortativity value indicates that individuals with similar collaboration behaviour are more likely to collaborate with each other.
@@ 4.1 Gender differences in career progression
To investigate gender differences in academic career progression, we examine three key aspects: the likelihood of becoming a PI, overall career length and career length after becoming a PI. We use the cumulative distribution function to show the probability of becoming a PI for each gender by a given time. We visualise the overall career length and career length after PI using the complementary cumulative distribution function for male and female authors. We can assess whether there are profound disparities in career progression between male and female authors by comparing these curves.
@@
Figure 5: Gender differences in career progression. These graphs show cumulative probability of being a PI and complementary cumulative probability during career length and career length after PI. Solid coloured lines represent cumulative and complementary cumulative density functions.
@@ 4.2 Assortativity in career progression
Figure 6 shows the assortativity of time spent at different career stages between co-authors. Overall we see slight positive assortativity in the career progression statistics. There does not appear to be any clear pattern in relation to the number of co-authored publications. These results suggest that frequent collaboration has little or no effect on the similarity of a co-author’s career progression. To validate the observed patterns in assortativity, we conducted a complementary analysis based on pairwise similarity in career outcomes between co-authors. As detailed in Appendix S7, the results confirm that higher collaboration frequency does not necessarily imply similar career progression.
Figure 6: Assortativity of career progression statistics according to the number of co-authored publications. The heat map provides a visual representation of the assortativity of each of the career progression statistics (y-axis) pairs of co-authors. The colour of each cell represents the assortativity according to the number of co-authored publications (x-axis). The bar chart along the top shows the logarithmic frequency of the co-authorship pairs. A positive assortativity value indicates that individuals with similar career progression are more likely to collaborate with each other.
@@ 5 Relationship between collaboration behaviour and career progression
Now we explore the relationship between collaboration behaviour, career progression and whether or not this relationship varies by gender. We use logistic regression models to investigate the relationship between collaboration statistics and the likelihood of an individual becoming a Principal Investigator (PI) and Accelerated Failure Time Models (AFTs) to analyse the association between collaboration and the time spent at different career stages.
@@ 5.1 Likelihood of becoming a PI
We assess the relationship between each of the adjusted collaboration statistics (network size, tie strength, clustering coefficient and co-authors per publication) and the likelihood of becoming a PI using logistic regression models. We also include gender and the interaction between gender and the collaboration statistic as predictors. We model the probability of an individual $i$ succeeding to become a PI ($s_{i}=1$) as follows:
@@
where $g_{i}$ is a binary variable that denotes the gender of the $i$-th individual, where $g_{i}=1$ indicates female and $g_{i}=0$ indicates male. The variable $c_{i}$ corresponds to the value of the specific collaboration statistic for the $i$-th individual. The model includes $\beta_{0}$ as the intercept term, $\beta_{1}$ for the collaboration static effect, $\beta_{2}$ for the gender effect and $\beta_{3}$ for the interaction effect between gender and the collaboration statistic. We also construct two separate models, one for male and one for female focal authors, to analyse the relationship within the two gender groups by fixing $\beta_{2}=\beta_{3}=0$.
@@
Figure 7: The forest plot of adjusted collaboration statistics on likelihood of being a PI. We present three types of model for each type of collaboration statistic (network size, tie strength, clustering coefficient and co-authors per publication). The interaction model refers to the full model described in Eq. (1) and includes the interaction effect of gender with the collaboration statistic. The female and male models assess the effect of the respective collaboration statistic on the likelihood of being a principal investigator (PI) for the separate groups of female and male authors. The Number column shows the sample size of each model. The markers show the coefficient estimates, with error bars indicating the 95% confidence intervals (CIs) for each estimate.
Figure 7 is a forest plot illustrating the association between the adjusted collaboration statistics and the likelihood of becoming a Principal Investigator (PI). The plot includes three types of model (interaction, female and male) for each adjusted collaboration statistic (network size, tie strength, clustering coefficient and co-authors per publication). Each row represents the value of the relevant model coefficient given in Eq. (1). In the interaction models, we include the coefficient $\beta_{3}$ to test whether the effect of each collaboration statistic on the likelihood of becoming a PI differs between male and female authors. A non-significant interaction term ($\beta_{3}$) indicates that the effect does not significantly vary across genders. In this specification, $\beta_{1}$ represents the effect of the collaboration statistic for male authors (the reference group), while the effect for female authors is given by the sum $\beta_{1}+\beta_{3}$, as also shown in the separate male and female models.
Figure 7 shows that the interaction terms between gender and most collaboration statistics are not statistically significant, except for tie strength. For men, a weaker mean tie strength is significantly associated with a greater likelihood of becoming a PI. However, the smaller number of female authors makes it harder to assess the statistical significance between mean tie strength and the likelihood of being a PI for women. We also see that an increased network size is significantly associated with a higher likelihood of becoming a PI, while a lower clustering coefficient and number of co-authors per publication are significantly associated with a greater likelihood of becoming a PI.